find the partial differential equations arising from z=xy+f(x^2+y^2)


1
Expert's answer
2020-09-28T15:46:36-0400

"\displaystyle z = xy + f(x^2 + y^2)\\


\textsf{Consider} \hspace{0.1cm} u = f(x^2 + y^2)\\

u = f(v), \hspace{0.1cm} v = x^2 + y^2\\

\frac{\partial u}{\partial v} = f'(v)\\


\therefore \frac{\partial u}{\partial x} = \frac{\partial u}{\partial v} \times \frac{\partial v}{\partial x}\hspace{0.1cm} \& \hspace{0.1cm} \frac{\partial u}{\partial y} = \frac{\partial u}{\partial v} \times \frac{\partial v}{\partial y}\\


\frac{\partial u}{\partial v} = f'(x^2 + y^2) \\

\frac{\partial v}{\partial x} = 2x \hspace{0.1cm} \& \hspace{0.1cm} \frac{\partial v}{\partial y} = 2y\\

\Rightarrow \frac{\partial z}{\partial x} = y + 2xf'(x^2 + y^2) \hspace{0.1cm}\&\hspace{0.1cm}\\

\frac{\partial z}{\partial y} = x + 2yf'(x^2 + y^2)\\



y\frac{\partial z}{\partial x} = y^2 + 2xyf'(x^2 + y^2) \hspace{0.5cm} (\textit{1}) \\


x\frac{\partial z}{\partial y} = x^2 + 2xyf'(x^2 + y^2)\hspace{0.5cm}(\textit{2})\\

\textsf{Subtracting}\hspace{0.1cm} (\textit{2}) \hspace{0.1cm} \textsf{from}\hspace{0.1cm} (\textit{1}), \hspace{0.1cm} \textsf{we have;}\\
y\frac{\partial z}{\partial x} - x\frac{\partial z}{\partial y} = y^2 - x^2\\

\therefore y\frac{\partial z}{\partial x} - x\frac{\partial z}{\partial y} = y^2 - x^2\\
\textsf{is a partial differential}\\\textsf{equation arising from}\hspace{0.1cm}z"